If a 10.0 kg tank of propane is completely consumed, what total mass of product is produced?
1
Expert's answer
2020-06-18T13:57:26-0400

Ans.

The balanced combustion reaction for propane is

C3H8(g)+5O2(g)→3CO2(g)+4H2O(g).

So, 1 mole of propane produces 3 moles of CO2 and 4 moles of H2O

Mole number of 10 kg propane is (10x103)/44.1 = 226.75 moles (Considering molecular weight of propane =  44.1 gm)

So, 226.75 moles of propane will produce (3 x 226.75) moles of CO2 and (4 x 226.75) moles of H2O

1 mole of CO2 = 44 gm CO2

So, (3 x 226.75) moles of CO2 = (44x3x226.75) gm CO2 = 29.931 kg CO2

 1 mole of H2O = 18 gm H2O

So, (4 x 226.75) moles of H2O = (18x4x226.75) gm CO2 = 16.321 kg H2O

Total mass of product produced = (29.931 + 16.321) kg = 46.252 kg (Ans.)